OpenStudy (anonymous):

v(x,y)=(y^2cos(xy^2)+ye^x, 2xycos(xy^2)+e^x) Find a function f such that v=grad f

OpenStudy (loser66):

is it not take integral of v?

OpenStudy (loser66):

@cinar

OpenStudy (loser66):

@Zizik v = grad f, it means v = gradient f, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

v(x,y) = (partial derivative of x, partial derivative of y)

OpenStudy (anonymous):

yes, taking integrals.

OpenStudy (anonymous):

\[f=\sin(xy^2)+ye^x\]

OpenStudy (anonymous):

\[\int\limits_{}^{}v(x,y)dx=\sin(xy^2)+ye^x+c_{1}(y)\\\ \int\limits_{}^{}v(x,y)dy=\sin(xy^2)+c_{1}(x)\]